Use the information in the table below to calculate the following ratios for Windswept Woodworks for year 1 and year 2. (Rouna. Interest coverage ratio (Assume that year 1 EBIT was $1,297 and year 1 interest expense was $120.) Year 2 interest coveragc. Current ratio Year 2 current ratio Year 1 current ratio d. Quick ratio Year 2 quick ratio Year 1 quick ratio
Answer #1

Answer:

a)
Interest Coverage Ratio = EBIT / Interest Expenses

Year 1:
Interest Coverage Ratio = 1,297 / 120
Interest Coverage Ratio = 10.81

Year 2:
EBIT = Sales – Cost of Goods Sold – Depreciation Expense
Depreciation = Accumulated Depreciation, Year 2 – Accumulated Depreciation, Year 1
Depreciation = $6782- $6662
Depreciation = $120

EBIT = $3018- $1500- $120
EBIT = $1398

Interest Coverage Ratio = 1398 / 140
Interest Coverage Ratio = 9.99

b)
Average Collection Period = 365 * Average Accounts Receivable / Sales

Year 1:
Average Accounts Receivable = (950 + 860) / 2
Average Accounts Receivable = $905

Average Collection Period = 365* 905 / 2,700
Average Collection Period = 122.34 days or 122 days

Year 2:
Average Accounts Receivable = (860 + 1316) / 2
Average Accounts Receivable = $1088

Average Collection Period = 365* 1088/ 3018
Average Collection Period = 131.58 days or 132 days

c)
Current Ratio = Current Assets / Current Liabilities

Year 1:
Current Assets = Accounts Receivable + Cash and Equivalents + Inventory
Current Assets = $860 + $158 + $1056
Current Assets = $2074

Current Liabilities = Accounts Payable + Other Current Liabilities
Current Liabilities = $414 + $96
Current Liabilities = $510

Current Ratio = 2074 / 510
Current Ratio = 4.07: 1

Year 2:
Current Assets = Accounts Receivable + Cash and Equivalents + Inventory
Current Assets = $1316 + $260 + $1050
Current Assets = $2626

Current Liabilities = Accounts Payable + Other Current Liabilities
Current Liabilities = $472 + $116
Current Liabilities = $588

Current Ratio = 2626 / 588
Current Ratio = 4.47: 1

d)
Quick Ratio = (Current Assets – Inventory) / Current Liabilities

Year 1:
Quick Ratio = (2074 – 1056) / 510
Quick Ratio = 1.996: 1

Year 2:
Quick Ratio = (2626– 1050) / 588
Quick Ratio = 2.68: 1